a bullet flying hirizontally can pierce 9 blocks of same thickness losing 75% of its total kinetic energy. the additonal number of blocks required to just arrest the bullet is:

Grade:12

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
14 years ago

Dear Ram Sai Charan,

Ans:- Let the initial velocity is U the final velocity is V Then from the given relation we have V²=0.75U²

or V=√3/2U

Let the retardation for each block is =f

then V²=U² - 2f(9x)

where x=thickness of each plate

So U²/4=18fx

Let the total no of block required to stop it from it's initial velocity is n

Then U²=2fnx

or   72fx=2fnx

or   n=36

Hence the excess no of plates required=36-9=27(ans) 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Ram Sai Charan !!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur